# Impact of the pre-equilibrium phase for the determination of nuclear geometry in high-energy isobar collisions

### v2

@@ Abstract
Ultrarelativistic isobar collisions have been proposed as a useful tool to investigate nuclear structure. These systems are not created in equilibrium, rather undergo a prethermalization phase. In this phase, some of the initial structure information may be lost and additional effects introduced. The objective of this paper is to study this possibility in the extreme case of a “free-streaming” pre-equilibrium phase. We do this by computing estimators for ratios of various measured (or measurable) quantities (elliptic and triangular flows, mean transverse momentum and associated cumulants, correlators between elliptic or triangular flows and mean transverse momentum, symmetric cumulant and two-plane correlator) and study their sensitivity to the duration of the free-streaming phase. We find that the correlators between elliptic or triangular flows and mean transverse momentum, the so-called $\rho_{2}$ and $\rho_{3}$, are indeed sensitive to the duration of the free-streaming phase and that the normalized symmetric cumulant, $\epsilon NSC(2,3)$ might also depend on this duration.
@@ I Introduction
The successful application of the hydrodynamic framework to describe high energy heavy-ion collisions has opened a new door to study nuclear structure. Experimental data from the collaborations STAR Adamczyk et al. (2015); Abdallah et al. (2022); Xu (2022), ALICE Acharya et al. (2018, 2022), CMS Sirunyan et al. (2019) and ATLAS Aad et al. (2020, 2023) as well as theoretical studies Heinz and Kuhlman (2005); Shou et al. (2015); Xu et al. (2018); Li et al. (2020); Giacalone et al. (2021a, b); Jia (2022a, b); Jia et al. (2022); Zhang and Jia (2022); Jia and Zhang (2023); Xu et al. (2021a); Bally et al. (2022a); Nijs and van der Schee (2023); Zhao et al. (2023a), have indicated the importance of nuclear deformations and radial distributions of protons and neutrons in determining the measured multiplicity distribution and collective flow in these collisions.
@@
Among deformed nuclei, the isobars ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r collisions at $\sqrt{s_{NN}}=200$ GeV studied by the STAR Collaboration Abdallah et al. (2022); Xu (2022) are of particular interest due to the similarities between the nuclei and the precision of the measurements. In these collisions, ${}^{96}_{44}$u exhibits higher particle production (by up to 10 percent), larger elliptic flow (by up to 3 percent), and smaller triangular flow (by 7 percent for central collisions) compared to ${}^{96}_{40}$r . These effects are relatively small, but still significant enough to be detected with high statistical precision and systematic uncertainties cancellation in ratios.
@@
On the theory side, a large number of collision simulations is needed to achieve such high precision, and the process must be repeated when poorly understood nuclear geometry parameters are changed. Hydrodynamic-based simulations are computationally expensive, so one approach is to use the known mapping between initial conditions and anisotropic flow Gardim et al. (2012a, 2015a); Niemi et al. (2013, 2016); Fu (2015). Numerous studies have employed the aforementioned methodology, such as Jia (2022a, b); Jia et al. (2022); Zhang and Jia (2022). Nevertheless, a limited number of works have utilized hydrodynamic simulations Xu et al. (2021a); Nijs and van der Schee (2023); Zhao et al. (2023a).
@@
State-of-the-art simulations of heavy-ion collisions are commonly based on a hybrid-chain of models, comprising four distinct stages: firstly, the generation of initial conditions; secondly, a prethermalization phase which serves to connect the initial conditions to the onset of hydrodynamic expansion; thirdly, the calculation of fluid expansion through the use of viscous relativistic hydrodynamics, and finally, a hadronic transport phase which occurs at the end of the hydrodynamic expansion. While the multiplicity distribution and collective flow are sensitive to the value of viscosity, as well as the inclusion of final state interactions and free streaming, it is expected that differences in the ratios of these observables only arise due to differences in nuclear structure. Although this assumption is reasonable, the precision of experimental data requires that it be tested. This has not be done systematically so far. An indirect study of viscosity and hadronic cascade effects using a transport model was recently conducted in Zhang et al. (2022).
@@
A fundamental physical question in the preceding hydrodynamic description is the time after which the viscous hydrodynamic description becomes valid. Early investigations indicated that to reproduce the observed large hadron momentum anisotropy, this time must be smaller than about 1 fm Heinz and Kolb (2002). However, studies in the strong and weak coupling limits have not fully explained this rapid thermalization, which triggered new approaches to understand when hydrodynamics does apply (see e.g. the reviews Romatschke (2017); Florkowski et al. (2018); Romatschke and Romatschke (2019)). The objective of this paper is to investigate the effect of the prethermalization on the description of isobar collisions. This is done in the extreme scenario where immediately after production, partons are free streamed until the beginning of the hydrodynamic evolution. We note that such an approach has been used in various recent Bayesian analysis, where the most likely range of parameter values for the hydrodynamic scenario is sought Bernhard et al. (2019); Everett et al. (2021a, b); Parkkila et al. (2021); Nijs et al. (2021a, b); Liyanage et al. (2023). We perform our study using well known estimators for elliptic and triangular flows, but we also present estimators that have received little attention for isobars for quantities such as mean transverse momentum and associated cumulants, correlators between elliptic/triangular flow and mean transverse momentum, symmetric cumulant, mixed harmonic correlator.
@@ II Setup of the simulations
Albeit many sophisticated models for nuclei exist from low-energy studies, the most popular way to model the nucleus in a ultra-relativistic collision is by sampling a Woods-Saxon distribution. The original distribution is spherically symmetric, hence not suitable for studying nuclei which are expected to have a deformed shape, such as ${}^{96}_{40}$r and ${}^{96}_{44}$u . The non-spherical shape is then taken into account by introducing an angular dependence in the nuclei radius, such as done in Jia (2022a); Giacalone et al. (2021b); Bally et al. (2022b). Hence, the probability $P(r,\,\theta,\,\varphi)$ of finding a nucleon in the position $\vec{r}=(r,\,\theta,\,\varphi)$ is proportional to the nuclear density profile
@@
where the nuclear radius $R_{0}$, diffuseness $a$, quadrupole deformation parameter $\beta_{2}$, triaxiality deformation parameter $\gamma$ and octupole deformation parameter $\beta_{3}$ characterize the nuclei structure. It is expected that variations in these parameters are reflected in final-state observables. Here, motivated by the values for these parameters present in KIBÉDI and SPEAR (2002); Pritychenko et al. (2016); Zhang and Jia (2022) for ${}^{96}_{44}$u and ${}^{96}_{40}$r , we sample 6 distinct cases, where we progressively tune our parameters from a triaxial ${}^{96}_{44}$u nucleus (case 1) to the ${}^{96}_{40}$r nucleus (case 6). The parameters used are shown in Table 1
{tblr}
colspec=—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—Q[c,50pt]—, hlines = black, .5pt, vlines = black, .5pt Case Label & $\mathbf{R_{0}}$ (fm) $\mathbf{a}$ (fm) $\mathbf{\beta_{2}}$ $\mathbf{\beta_{3}}$ $\mathbf{\gamma}$
Case 1 5.09 0.46 0.16 0 $\pi/6$ [Uncaptioned image]
Case 2 5.09 0.46 0.16 0 0 [Uncaptioned image]
Case 3 5.09 0.46 0.16 0.20 0 [Uncaptioned image]
Case 4 5.09 0.46 0.06 0.20 0 [Uncaptioned image]
Case 5 5.09 0.52 0.06 0.20 0 [Uncaptioned image]
Case 6 5.02 0.52 0.06 0.20 0 [Uncaptioned image]
@@
In Table 1, when going from one line to the next, only one parameter is changed. Cases 1 and 2 are similar except that the former is a triaxial ellipsoid and the later is a spheroid. An octupole deformation (pear shape) has been introduced when going from case 2 to 3. When going from case 3 to 4, the quadrupole deformation (ellipsoid shape) is reduced. Cases 4 and 5 are similar except that the later has a larger diffuseness. Lastly, the nuclear radius is reduced when going from case 5 to 6. These systematic modifications in the nuclear structure parameters allow to investigate eventual changes in the results when gradually moving from what is typically considered a Ruthenium nucleus (case 2) to what typically is used as a Zirconium nucleus (case 6). Effects related to “neutron-skin” have not been included in our simulations, therefore, protons and neutrons follow the same nuclear distribution. The reader may refer to Xu et al. (2021b) for a study on the neutron-skin effects.
@@
Due to its flexibility and the ability to accept pre-generated (possibly deformed) nuclear configurations for projectile and target, the TRENTo model Moreland et al. (2015) has been used to generate initial conditions for our simulations. Two nuclear configurations (1) are randomly selected from a set of 50.000 pre-generated options and subsequently randomized in orientation. The Monte-Carlo Glauber model, where a value for the nucleon-nucleon cross-section $\sigma_{\rm NN}$ is needed, is then applied to identify which nucleons will participate in the collision, and to each participating nucleon, a density function $\rho_{i}(x,y,z)$ is assigned, namely a Gaussian function with standard deviation given by the nucleon width $w$. Each nucleon density is allowed to fluctuate according to a Gamma distribution with shape parameter $k$. By summing these densities and integrating over the $z$ direction for each nucleus, the participant thickness $T_{A,B}$ is obtained. These thicknesses are then used to calculate the energy density $\epsilon(x,y,\eta_{s}=0)$ as a reduced thickness between the two participant thicknesses given by
$\lim_{\tau\to 0^{+}}\tau\epsilon(x,y,\eta_{s}=0)=N\left[\frac{T_{A}(x,y)+T_{B}(x,y)}{2}\right]^{1/p}\,,$ (2)
@@
where $N$ is a normalization constant and $p$ is called the reduced thickness parameter. Since in this work we will be dealing solely with ratios of quantities such as eccentricities and average radius of the system, the exact value of the normalization should not matter. However, in order to actually produce initial conditions with TRENTo , one needs to specify a value for its physical parameters ($N$, $w$, $\sigma_{\rm NN}$, $k$, $p$). In our simulations, these parameters were chosen to be the same as in Everett et al. (2021b) (because their analysis was done for both RHIC and LHC energies) and are outlined in Table 2 for completeness.
@@
Table 2: TRENTo parameters used to generate initial conditionsEverett et al. (2021b).
@@
After the collision takes place, one needs to specify how long the produced system will stay in the prethermalization phase. Given the model chosen in this work, such characteristic is controlled by the free-streaming time, $\tau_{FS}$. It is worth noting that at the moment all assumptions related to the modeling of pre-equilibrium phase are ad-hoc and such a quantity is currently chosen to have either a constant value or is allowed to carry a dependence with the collision centrality.
@@
Regardless of the case, the value of $\tau_{FS}$ or its behavior with the collision centrality started to be determined (together with many other parameters needed in a complete hybrid hydrodynamic simulation) via Bayesian inference techniques Moreland et al. (2020); Bernhard et al. (2019); Everett et al. (2021b, a); Nijs et al. (2021a); Parkkila et al. (2021); Nijs and van der Schee (2022); Parkkila et al. (2022). This does not mean, however, that the main characteristics of this phase are well understood. In fact, it depends on the many particularities of each statistical analysis as different collaborations have obtained values for $\tau_{FS}$ that vary more than 50%: for instance, Parkkila et al. (2021) extracted $\tau_{FS}=0.901$ fm/c by considering Pb+Pb collisions at 5.02 TeV; on the other hand, Parkkila et al. (2022) reported $\tau_{FS}=0.71$ fm/c improving on the previous study by considering a simultaneous analysis of Pb+Pb data at 2.76 TeV and 5.02 TeV. When allowing the free-streaming to change with centrality as in Everett et al. (2021b), with parameters from the Grad viscous correction model, for Au+Au collisions at 200 GeV, $1.36\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.51$, a variation of almost 50% when compared to the improved estimate from Parkkila et al. (2022). Lastly, a simultaneous analysis of p+Pb and Pb+Pb data at 5.02 TeV presented in Moreland et al. (2020) yielded $\tau_{FS}=0.37$ fm/c. While a smaller free-streaming time is expected for p+A collisions due to the system size after the collision takes place, this estimate (which is also applied to Pb+Pb collisions if one makes use of the model outlined in Moreland et al. (2020)) is $\sim 0.59\%$ smaller than the value quoted in Parkkila et al. (2021) which only considered Pb+Pb data at the same collision energy.
@@
Given the current lack of a detailed understanding of the prethermalization phase and the large uncertainty in the extracted values for $\tau_{FS}$, in the present study we consider three different choices for the free-streaming time: (i) $\tau_{FS}=0.4$ fm/c, (ii) $\tau_{FS}=1.0$ fm/c and (iii) a centrality-dependent free-streaming time as implemented in Everett et al. (2021b): $\tau_{FS}=\tau_{R}(\{\epsilon\}/\{\epsilon_{R}\})^{\alpha}$, where $\{\epsilon\}$ denotes a (particular) average initial energy density in the transverse plane for a given collision, $\{\epsilon_{R}\}=4.0$ GeV/fm2, $\tau_{R}=1.46$ fm/c and $\alpha=0.031$ are constants that were adjusted by comparison with experimental data. These choices cover the range of values determined by recent Bayesian analysis and allow for a systematic investigation of eventual impacts that the prethermalization phase may have on quantities related to the system shape. In Fig. 1, the free-streaming time for the centrality-dependent case is shown as function of $\{\epsilon\}$ for Ru (very similar results hold for Zr). Centrality windows were estimated from ordering the total energy $E$ (integral on the transverse plane of the energy density in eq. (2) of 1 million TRENTo initial conditions in descending order for each colliding system. One sees that the free-streaming time varies between $\approx$ 1.35 and $\approx$ 1.49 fm in such a case.
@@
Figure 1: Free-streaming time as function of $\{\epsilon\}$ in a mid-rapidity slice with $\tau_{R}=1.46$ fm/c, $\{\epsilon_{R}\}=4$ GeV/fm2 and $\alpha=0.031$. Centralities windows are also shown (see text). Very similar results are obtained for ${{}_{40}^{96}Zr}+{{}_{40}^{96}Zr}$.

### v4

# Impact of the pre-equilibrium stage for the determination of nuclear geometry in high-energy isobar collisions

@@ Abstract
Ultrarelativistic isobar collisions have been proposed as a useful tool to investigate nuclear structure. These systems are not created in equilibrium, rather undergo a pre-thermalization stage. In this stage, some of the initial structure information may be lost and additional effects introduced. The objective of this paper is to study this possibility in the extreme case of a “free-streaming” pre-equilibrium stage. We do this by computing estimators for ratios of various measured (or measurable) quantities (elliptic and triangular flows, mean transverse momentum and associated cumulants, correlators between elliptic or triangular flows and mean transverse momentum, symmetric cumulant and two-plane correlator) and study their sensitivity to the duration of the free-streaming stage. We find that the correlators between elliptic or triangular flows and mean transverse momentum, the so-called $\rho_{2}$ and $\rho_{3}$, are indeed sensitive to the duration of the free-streaming stage and that the normalized symmetric cumulant, $\varepsilon NSC(2,3)$ might also depend on this duration.
@@ I Introduction
The successful application of the hydrodynamic framework to describe high energy heavy-ion collisions has opened a new door to study nuclear structure. Experimental data from the collaborations STAR [1, 2, 3], ALICE [4, 5], CMS [6] and ATLAS [7, 8] as well as theoretical studies [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], have indicated the importance of nuclear deformations and radial distributions of protons and neutrons in determining the measured multiplicity distribution and collective flow in these collisions.
@@
Among deformed nuclei, the isobars ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r collisions at $\sqrt{s_{NN}}=200$ GeV studied by the STAR Collaboration [2, 3] are of particular interest due to the similarities between the nuclei and the precision of the measurements. In these collisions, ${}^{96}_{44}$u exhibits higher particle production (by up to 10 percent), larger elliptic flow (by up to 3 percent), and smaller triangular flow (by 7 percent for central collisions) compared to ${}^{96}_{40}$r . These effects are relatively small, but still significant enough to be detected with high statistical precision and systematic uncertainties cancellation in ratios.
@@
On the theory side, a large number of collision simulations is needed to achieve such high precision, and the process must be repeated when poorly understood nuclear geometry parameters are changed. Hydrodynamic-based simulations are computationally expensive, so one approach is to use the known mapping between initial conditions and anisotropic flow [47, 24, 48, 25, 26]. Numerous studies have employed the aforementioned methodology, such as [15, 16, 17, 18]. Nevertheless, a limited number of works have utilized hydrodynamic simulations [20, 22, 23].
@@
State-of-the-art simulations of heavy-ion collisions are commonly based on a hybrid-chain of models, comprising four distinct stages: firstly, the generation of initial conditions; secondly, a pre-thermalization stage which serves to connect the initial conditions to the onset of hydrodynamic expansion; thirdly, the calculation of fluid expansion through the use of viscous relativistic hydrodynamics, and finally, a hadronic transport phase which occurs at the end of the hydrodynamic expansion. While the multiplicity distribution and collective flow are sensitive to the value of viscosity, as well as the inclusion of final state interactions and free streaming, it is expected that differences in the ratios of these observables only arise due to differences in nuclear structure. Although this assumption is reasonable, the precision of experimental data requires that it be tested. This has not be done systematically so far. An indirect study of viscosity and hadronic cascade effects using a transport model was recently conducted in [27].
@@
A fundamental physical question in the preceding hydrodynamic description is the time after which the viscous hydrodynamic description becomes valid. Early investigations indicated that to reproduce the observed large hadron momentum anisotropy, this time must be smaller than about 1 fm [28]. However, studies in the strong and weak coupling limits have not fully explained this rapid thermalization, which triggered new approaches to understand when hydrodynamics does apply (see e.g. the reviews [29, 30, 31]). The objective of this paper is to investigate the effect of the pre-thermalization on the description of isobar collisions. This is done in the extreme scenario where immediately after production, partons are free streamed until the beginning of the hydrodynamic evolution. We note that such an approach has been used in various recent Bayesian analysis, where the most likely range of parameter values for the hydrodynamic scenario is sought [32, 33, 34, 35, 36, 37, 38]. We perform our study using well known estimators for elliptic and triangular flows, but we also present estimators that have received little attention for isobars for quantities such as mean transverse momentum and associated cumulants, correlators between elliptic/triangular flow and mean transverse momentum, symmetric cumulant, mixed harmonic correlator.
@@ II Setup of the simulations
Albeit many sophisticated models for nuclei exist from low-energy studies, the most popular way to model the nucleus in a ultra-relativistic collision is by sampling a Woods-Saxon distribution. The original distribution is spherically symmetric, hence not suitable for studying nuclei which are expected to have a deformed shape, such as ${}^{96}_{40}$r and ${}^{96}_{44}$u . The non-spherical shape is then taken into account by introducing an angular dependence in the nuclei radius, such as done in [15, 14, 39]. Hence, the probability $P(r,\,\theta,\,\varphi)$ of finding a nucleon in the position $\vec{r}=(r,\,\theta,\,\varphi)$ is proportional to the nuclear density profile
@@
where the nuclear radius $R_{0}$, diffuseness $a$, quadrupole deformation parameter $\beta_{2}$, triaxiality deformation parameter $\gamma$ and octupole deformation parameter $\beta_{3}$ characterize the nuclei structure. It is expected that variations in these parameters are reflected in final-state observables. Here, motivated by the values for these parameters present in [40, 41, 18] for ${}^{96}_{44}$u and ${}^{96}_{40}$r , we sample 6 distinct configurations, where we progressively tune our parameters from a triaxial ${}^{96}_{44}$u nucleus (configuration 1) to the ${}^{96}_{40}$r nucleus (configuration 6). The parameters used are shown in Table 1.
@@
Configuration Label $\mathbf{R_{0}}$ (fm) $\mathbf{a}$ (fm) $\mathbf{\hskip 5.69046pt\beta_{2}\hskip 5.69046pt}$ $\mathbf{\hskip 5.69046pt\beta_{3}\hskip 5.69046pt}$ $\mathbf{\gamma}$
Configuration 1 5.09 0.46 0.16 0 $\pi/6$
Configuration 2 5.09 0.46 0.16 0 0
Configuration 3 5.09 0.46 0.16 0.20 0
Configuration 4 5.09 0.46 0.06 0.20 0
Configuration 5 5.09 0.52 0.06 0.20 0
Configuration 6 5.02 0.52 0.06 0.20 0
@@
In Table 1, when going from one line to the next, only one parameter is changed. Configurations 1 and 2 are similar except that the former is a triaxial ellipsoid and the later is a spheroid. An octupole deformation (pear shape) has been introduced when going from configuration 2 to 3. When going from case 3 to 4, the quadrupole deformation (ellipsoid shape) is reduced. Configurations 4 and 5 are similar except that the later has a larger diffuseness. Lastly, the nuclear radius is reduced when going from configuration 5 to 6. These systematic modifications in the nuclear structure parameters allow to investigate eventual changes in the results when gradually moving from what is typically considered a Ruthenium nucleus (configuration 2) to what typically is used as a Zirconium nucleus (configuration 6). The effect of these different parameters in the shape of the nucleus can be seen in Fig. 1. Effects related to “neutron-skin” have not been included in our simulations, therefore, protons and neutrons follow the same nuclear distribution. The reader may refer to [42] for a study on the neutron-skin effects.
Figure 1: Shapes of different configurations in the nuclei, using the parameters of Table 1 in Eq. (1).
@@
Due to its flexibility and the ability to accept pre-generated (possibly deformed) nuclear configurations for projectile and target, the TRENTo model [43] has been used to generate initial conditions for our simulations. Two nuclear configurations (1) are randomly selected from a set of $50000$ pre-generated options and subsequently randomized in orientation. The Monte-Carlo Glauber model, where a value for the nucleon-nucleon cross-section $\sigma_{\rm NN}$ is needed, is then applied to identify which nucleons will participate in the collision, and to each participating nucleon, a density function $\rho_{i}(x,y,z)$ is assigned, namely a Gaussian function with standard deviation given by the nucleon width $w$. Each nucleon density is allowed to fluctuate according to a Gamma distribution with shape parameter $k$. By summing these densities and integrating over the $z$ direction for each nucleus, the participant thickness $T_{A,B}$ is obtained. These thicknesses are then used to calculate the energy density $\epsilon(x,y,\eta_{s}=0)$ as a reduced thickness between the two participant thicknesses given by
$\bar{\epsilon}(x,y)\equiv\lim_{\tau\to 0^{+}}\tau\epsilon(x,y,\eta_{s}=0)=N\left[\frac{T_{A}^{p}(x,y)+T_{B}^{p}(x,y)}{2}\right]^{1/p}\,,$ (2)
@@
where $N$ is a normalization constant and $p$ is called the reduced thickness parameter. Since in this work we will be dealing solely with ratios of quantities such as eccentricities and average radius of the system, the exact value of the normalization should not matter. However, in order to actually produce initial conditions with TRENTo , one needs to specify a value for its physical parameters ($N$, $w$, $\sigma_{\rm NN}$, $k$, $p$). In our simulations, these parameters were chosen to be the same as in [34] (because their analysis was done for both RHIC and LHC energies) and are outlined in Table 2 for completeness.
Table 2: TRENTo parameters used to generate initial conditions[34]
@@
After the collision takes place, one needs to specify how long the produced system will stay in the pre-thermalization stage. Given the model chosen in this work, such characteristic is controlled by the free-streaming time, $\tau_{FS}$. It is worth noting that at the moment all assumptions related to the modeling of pre-equilibrium stage are ad-hoc and such a quantity is currently chosen to have either a constant value or is allowed to carry a dependence with the collision centrality.
@@
Regardless of the case, the value of $\tau_{FS}$ or its behavior with the collision centrality started to be determined (together with many other parameters needed in a complete hybrid hydrodynamic simulation) via Bayesian inference techniques [44, 32, 34, 33, 36, 35, 45, 46]. This does not mean, however, that the main characteristics of this stage are well understood. In fact, it depends on the many particularities of each statistical analysis as different collaborations have obtained values for $\tau_{FS}$ that vary more than 50%: for instance, [35] extracted $\tau_{FS}=0.901$ fm/c by considering Pb+Pb collisions at 5.02 TeV; on the other hand, [46] reported $\tau_{FS}=0.71$ fm/c improving on the previous study by considering a simultaneous analysis of Pb+Pb data at 2.76 TeV and 5.02 TeV. When allowing the free-streaming to change with centrality as in [34], with parameters from the Grad viscous correction model, for Au+Au collisions at 200 GeV, $1.36\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.51$, a variation of almost 50% when compared to the improved estimate from [46]. Lastly, a simultaneous analysis of p+Pb and Pb+Pb data at 5.02 TeV presented in [44] yielded $\tau_{FS}=0.37$ fm/c. While a smaller free-streaming time is expected for p+A collisions due to the system size after the collision takes place, this estimate (which is also applied to Pb+Pb collisions if one makes use of the model outlined in [44]) is $\approx 0.59\%$ smaller than the value quoted in [35] which only considered Pb+Pb data at the same collision energy.
@@
Given the current lack of a detailed understanding of the pre-thermalization stage and the large uncertainty in the extracted values for $\tau_{FS}$, in the present study we consider three different choices for the free-streaming time: (i) $\tau_{FS}=0.4$ fm/c, (ii) $\tau_{FS}=1.0$ fm/c and (iii) a centrality-dependent free-streaming time as implemented in [34]: $\tau_{FS}=\tau_{R}(\{\bar{\epsilon}\}/\bar{\epsilon}_{R})^{\alpha}$, where $\{{\bar{\epsilon}}\}$ denotes a (particular) average initial energy density in the transverse plane for a given collision, $\bar{\epsilon}_{R}=4.0$ GeV/fm2, $\tau_{R}=1.46$ fm/c and $\alpha=0.031$ are constants that were adjusted by comparison with experimental data. These choices cover the range of values determined by recent Bayesian analysis and allow for a systematic investigation of eventual impacts that the pre-thermalization stage may have on quantities related to the system shape. In Fig. 2, the free-streaming time for the centrality-dependent case is shown as function of $\{{\bar{\epsilon}}\}$ for Ru (very similar results hold for Zr). Centrality windows were estimated from ordering the total energy $E$ (integral on the transverse plane of the energy density in eq. (2), $E=N\int d^{2}x_{\perp}\left[\frac{T_{A}^{p}(x,y)+T_{B}^{p}(x,y)}{2}\right]^{1/p}$) of 1 million TRENTo initial conditions in descending order for each colliding system. One sees that the free-streaming time varies between $\approx$ 1.35 and $\approx$ 1.49 fm in such a case.
In Fig. 2, $\tau_{FS}$, the time when free-streaming stops and hydrodynamics is switched on, increases for larger energy densities. Though this is discussed and some arguments for this put forward in [34], one might expect the opposite: for larger densities, more collisions happen and hydrodynamics starts earlier. Therefore we have also included a curve that corresponds to a negative value for $\alpha$. For central collisions, which are usually considered more interesting for isobar collisions, both positive and negative $\alpha$ lead to similar fairly constant values. However we will discuss both cases when necessary below.
@@
Figure 2: Free-streaming time as function of $\{{\bar{\epsilon}}\}$ in a mid-rapidity slice with $\tau_{R}=1.46$ fm/c, $\bar{\epsilon}_{R}=4$ GeV/fm2 and $\alpha=0.031$ [34]. Centrality windows are also shown (see text). Very similar results are obtained for ${{}_{40}^{96}Zr}+{{}_{40}^{96}Zr}$. The $\alpha=-0.031$ case is shown by the red dot-dashed line. Both positive and negative $\alpha$’s lead to comparable values of $\tau_{FS}$ for central collisions.